OpenStudy (anonymous):

(cosx-sinx)^2+(cosx+sinx)^2=2 Please help will give medal

OpenStudy (anonymous):

Prove the identity

OpenStudy (anonymous):

Will give medal

OpenStudy (math&ing001):

Distribute (cosx-sinx)^2 and (cosx+sinx)^2 then use the property cos(x)^2+sin(x)^2=1

OpenStudy (tkhunny):

Have you considered squaring the indicated values?

OpenStudy (anonymous):

How would I square it

OpenStudy (alekos):

(cosx - sinx)^2 = cos^2x -2cosxsinx +sin^2x (cosx+sinx)^2 = cos^2x+2cosxsinx+sin^2x then add them together

OpenStudy (tkhunny):

OR - Save some headaches with a little more thought. Substitute \(\cos(x) + \sin(x) = \sqrt{2}\sin(x+\pi/4)\) Substitute \(\cos(x) - \sin(x) = \sqrt{2}\cos(x+\pi/4)\) Rather pops right out after that.

OpenStudy (tkhunny):

You should know how to square a binomial. You WILL be required to have this skill on examination. \((a+b)^{2} = a^{2} + 2ab + b^{2}\) You've done it lots of time, haven't you?

OpenStudy (anonymous):

2cos pi/4

OpenStudy (anonymous):

No I suck at trig

OpenStudy (tkhunny):

What is that? It bears no resemblance to anything else we've mentioned.

OpenStudy (alekos):

yeah, i'll say

OpenStudy (tkhunny):

Ah. Well, it is time to stop sucking. Besides, squaring a binomial is from Algebra I. Come one. Make a better effort. Slowly. Patiently. You'll get it.

OpenStudy (anonymous):

I'll do my best

OpenStudy (anonymous):

can some show me how to do this

OpenStudy (alekos):

i just gave you half the answer. just add the two expressions

OpenStudy (anonymous):

cosxsinx

OpenStudy (anonymous):

is that right